OpenStudy (jamanorthi):

In the expansion of (2x^2-3x-1) (x+5) what is the coefficient of x^2?

OpenStudy (anonymous):

Consider which terms contribute to an \(x^2\) term in the product. \(2x^2\times5=10x^2\) and \(-3x\times x=-3x^2\)